1.    Consider an ordinary shower where hot water at 60°C is mixed with cold water at 10°C. If it is desired that a steady stream of warm water at 45°C be supplied, determine the ratio of the mass flow rates of the hot to cold water. Assume the heat losses from the mixing chamber to be negligible and the mixing to take place at a pressure of 150 kPa. 


1
Expert's answer
2022-02-14T00:07:01-0500

Let "M_h=" flow rate of hot water

"M_c" "=" flow rate of cold water



Energy lost by hot water"=" Energy gained by cold water


"M_Hc(60-45)=M_cc(45-10)"


Where c is specific heat capacity of water


"\frac{M_H}{M_c}=\frac{45-10}{60-45}=\frac{35}{15}=\frac{7}{3}"



"M_H:M_c=7:3"